An automatic coin counter at a bank indicates that $65 in dimes and quarters was processed. If there were 389 coins how many of each were there?
Answer by checkley71(8403)

D+Q=389 OR D=389-Q
10(389-Q)+25Q=6500
3890-10Q+25Q=6500
15Q=6500-3890
15Q=2610
Q=2610/15
Q=174 QUARTERS
389-174=215 DIMES.
PROOF
10*215+25*174=6500
2150+4350=6500
6500=6500